OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5 a 2 b -2 c 1 d -1

OpenStudy (texaschic101):

substitution method will work just fine.... x - 2y = - 5 ---> x = 2y - 5 now sub 2y - 5 in for x in the 1st equation 2x + 3y = 4 2(2y - 5) + 3y = 4 now distribute the 2 through the parenthesis. Combine like terms. Get the y's on one side and everything else on the other side. Solve for y. Once you have y solved, sub that answer into either of the original equations to find x. Can you take it from here ?

OpenStudy (raden_zaikaria):